Algebra

The chance of one event happening is the square of the chance of second event ,but odds against the first are the cubes of the odds against the second. Find the chances of each.( assume that both events are neither sure nor impossible)

utkarsha  sonje
14 Years agoGrade

2 Answers

Chetan Mandayam Nayakar
14 Years ago

let x1 and x2 be the chances

x1=x2^2

1-x1=(1-x2)^3

solve for x1 and x2

 

saurabh
8 Years ago
let x be the occurrence of first event and ybe the occurrence of second event. 
according to question x=y2
and but also given (1-X)=(1-y)3 solve the both  equations ,we get y=1/3 and x=1/9.